OpenStudy (anonymous):

f(x)= ((x-4)(x+5)/((13x+4)(4-x)) What graphical feature occurs at x=4? (select all that apply) hole intercept vertical asymptote

OpenStudy (jkristia):

\[f(x) = \frac{(x-4)(x+5)}{(x-4)(13x+4)}\] This function contains a removeable discontinuity. Obviously if you set x = 4 you get 0/0 which is undefined but since you have (x-4) in both numerator and denominator you can remove those (they cancel out), just remember that the function is undefined for x = 4. So your f(x) when x is really close to 4 can be calculates as if x is = 4 (except it cannot be exactly 4) \[f(x) = \lim_{x \rightarrow 4} \frac{(x+5)}{(13x+4)} = \frac{9}{56}\] so f(x) is 9/56 when x is really really close to 4, so what is the behavior when x is 4 ?

OpenStudy (anonymous):

hole

OpenStudy (jkristia):

yep

OpenStudy (anonymous):

Just a minor thing but the lim should be \[-\frac{ 9 }{56 }\] if the numerator contains (x-4) and the denominator contains (4-x).

OpenStudy (jkristia):

oops - sorry, my mistake, I didnt notice different terms. As you can see I have both numerator and denominator as (x-4). But you are right, one needs to be multiplied by -1. Thank you for pointing this out.

OpenStudy (anonymous):

Just a minor thing but the lim should be